A wire loop of resistance 10Ω and radius 10cm is kept in the plane of this paper in a uniform magnetic field B. The direction of B is perpendicular to the plane of the page and points out of it and its magnitude is increasing at the rate of 0.50T/s. Determine the magnitude and direction of the induced current in the loop.
1
Expert's answer
2020-03-09T11:19:17-0400


From the law of electromagnetic induction we write


"\Epsilon_i=\frac{\Delta\Phi}{\Delta t}=\frac{\Delta(B \cdot S)}{\Delta t}=(\frac{\Delta B }{\Delta t})\cdot S"


Where

"S=\pi \cdot r^2"


"(\frac{\Delta B }{\Delta t})" -The rate of change of the magnetic field

Then the current is

"I=\frac{\Epsilon_i}{R}=\frac{(\frac{\Delta B }{\Delta t})\cdot S}{R}=\frac{(\frac{\Delta B }{\Delta t})\cdot \pi \cdot r^2}{R}=\frac{(0.5)\cdot 3.14 \cdot 0.1^2}{10}=1.57 mA"